OpenStudy (anonymous):

Solve the equation for a 3a^2+22a+35=0

OpenStudy (callisto):

\[3a^2+22a+35=0\]\[(3a+7)(a+5)=0\]3a+7 =0 or a+5=0 Can you do it from here?

OpenStudy (anonymous):

I was thinking about using this quadratic formular